# Efficient Measure of Information Backflow with Quasi-stochastic Process

### v1

@@ Abstract
where $j\equiv(j_{1},j_{2})\in\mathbb{Z}_{2}\times\mathbb{Z}_{2}$, and $\sigma_{x},\sigma_{y},\sigma_{z}$ are qubit Pauli operators. The construction for discrete Wigner representation is more involved and since we are only working with qutrit system in the example below we show the frame operators for $d=3$ in the Appendix.
We ascertain that the results below are frame independent by repeating the calculation with other representation, namely, the symmetric informationally-complete (SIC)-POVM representation [36].
Example 1: Pure decoherence model—Let us start with the prototypical model of a two-level system in a bosonic bath environment undergoing a pure decoherence dynamics [11]. The evolution of the system’s state is given by the master equation
@@
where the time-dependent decay rate $\gamma(t)=-\frac{1}{G(t)}\frac{d}{dt}G(t)$ with $G(t)$ being the decoherence function. From this, the dynamical map reads
@@
Note that the map $\Lambda_{t}$ is CPTP for $G(t)\leq 1$.
To find the non-Markovianity criteria, we first find the quasi-stochastic representation of $\Lambda_{t}$ following the Wootters’ representation:
@@
where $G\equiv G(t)$ for simplicity. Notice that here $S^{\Lambda_{t}}$ is self-transpose, hence $\Lambda_{t}$ is its self-dual, i.e., $\Lambda_{t}^{\dagger}=\Lambda_{t}$. More importantly, the quasi-bistochastic map is represented non-negatively (same with SIC-POVM representation) that affords us to use the main result in Eq. 9. In fact, all the examples in this paper have a map that are non-negative. Calculating the LHS of Eq. (11), we find its eigenvalues are given by $\{1,1,G^{2},G^{2}\}$. Using Eq. (Efficient Measure of Information Backflow with Quasi-stochastic Process), we obtain that
@@
[…]
Example 2: Dissipation model—Next, we consider the same system and bath environment but with system-environment interaction that allows for dissipation in the populations too. This evolution is governed by the time-local master equation
@@
$\Lambda_{t}:\rho(0)\rightarrow\rho(t)=\begin{bmatrix}\rho_{00}+(1-|G(t)|^{2})\rho_{11}&G(t)^{*}\rho_{01}\\ G(t)\rho_{10}&|G(t)|^{2}\rho_{11}\end{bmatrix}\,.$ (19)
@@
$\zeta(t)=8|G(t)|^{3}\frac{d}{dt}|G(t)|=-4\gamma(t)|G(t)|^{4}\,,$ (20)
and identify that non-Markovianity is observed when $\gamma(t)<0$. Again, the witness here coincides with the BLP and CP-divisibility measure [1].
Example 3: Random unitary channel—In the last example, we shall look into evolution under random unitary channel of a qubit and qutrit system. The general qudit system has been studied in [37]. Here, we consider the master equation
@@
[…]
For $d=2$, the Weyl operators reduce to the identity and Pauli operators. In this case, we find the criterion for Markovian condition to be
@@
This coincides with the BLP and P-divisibility condition as expected [38] but differs from the CP-divisibility condition, which is $\gamma_{k}(t)\geq 0$. This is an instance of the departure of the Markovian condition between information backflow and divisibility with the former being weaker.
In $d=3$, we find the Markovian criterion as
@@
To our knowledge, characterization based on BLP and P-divisibility appears to be highly non-trivial and only sufficient (but not necessary) condition had been obtained:
@@
for all triples $\{i_{1},i_{2},i_{3}\}\subset\{1,\dots,8\}$. Similar to the qubit case, the evolution is CP-divisible if all rates are non-negative, $\gamma_{\alpha}(t)\geq 0$.
@@
Discussions—The monotonic relation in Eq. (11) provides an interesting perspective on non-Markovianity between classical and quantum theory. In the classical case, evolution is generally described by a stochastic map. Bistochastic map, a subset of all stochastic maps, has a special property such that its transpose returns a valid map as well. One can think of the transpose as an attempt to ‘reverse’ the transition probabilities between the input and output. Recent works [39, 40, 41] have also pointed out that Hermitian adjoint and transpose of a quantum and classical process, respectively, correspond to the time-reversal of the process. Interestingly, under QPR the time-reversal notion has been unified for both theories with the transpose operation. Reversal of the process, however, does not always help in reversing the state of the system. This is also reflected in the monotonic relation, in both forms of $S^{T}S$ and $SS^{T}$, that it will mix any probability distribution further. In contrast, applying forward and reverse quantum processes successively, or vice versa, on any state may result in a momentary revival of quantum features albeit still leading to noise eventually.
@@
It should be noted that ‘reversal’ is not equivalent to the inverse, which is generally not a valid map except for unitaries. [^2] For classical theory, unitary processes are described by permutation matrices — matrices with a single $1$ in each row and columns and 00 elsewhere Nevertheless, for an irreversible process one can still construct a map that approximately invert the forward process known as Bayesian inference. Remarkably, the canonical quantum Bayesian inference, Petz map, also contains $S^{T}$ which might imply that it is a useful ingredient in retrieving loss information [42].
@@
In formulating the condition for Markovianity, it is also required that quantum maps are represented non-negatively. This is seemingly counterintuitive to current perception where it is widely believed that the appearance of negativity is necessary for a lot of quantum phenomena. Thus, one would also think that quantum non-Markovianity, a non-classical phenomenon, would occur due to negativity in the dynamical map. It would be interesting to show if a negative quasi-stochastic map corresponding to a Markovian evolution would still satisfy the monotonic relation in Eq. (9). Alternatively, it would also be interesting if it is always possible to find a frame representation that will return a non-negative quasi-stochastic matrix.
@@
Although the QPR of the dynamical maps studied above are represented non-negatively, we would like to point out that the generator $\mathcal{L}$ has a non-classical representation. More explicitly, we find $L=[L_{jk}]$, $L_{jk}=\text{Tr}\left[F_{j}\mathcal{L}[G_{k}]\right]\in\mathbb{R}$, which is a $d^{2}\times d^{2}$ matrix that satisfies $\sum_{j}L_{jk}=0\,\forall\,k$. This is the quantum analogue of the classical Kolmogorov generator [43] that additionally satisfies $L_{jk}\geq 0$ for $j\neq k$, i.e., its off-diagonal elements are non-negative. For the pure decoherence and random unitary channel models, we observe that the off-diagonal entries of $L$ become negative when non-Markovianity condition is satisfied and non-negative otherwise. However, for dissipative model the matrix $L$ has negative non-diagonal entries in general. From here, it suggests that negativity in the generator appears to be necessary but not sufficient for non-Markovianity. This observation has also been pointed out in [44], where they studied its connection to classical and quantum P-divisibility. It remains unclear how exactly negativity in the generator of quantum evolution contributes to the degree of non-Markovianity. We believe that the study using QPR would be crucial in gaining a fundamental understanding of the difference between classical and quantum processes.
@@
Conclusions – In this Letter, we have proposed a new witness and measure of non-Markovianity using the formalism of quasiprobability representation. In particular, the measure detects information backflow through the monotonic relation on the product of quasi-bistochastic map, representing the dynamical map, and its transpose. The measure is explicitly dependent only on the dynamical map, thus foregoing any optimization task that plagues much of the previously proposed measure. This is particularly useful when studying evolution of a quantum system of large dimensions. We illustrate the use over several paradigmatic examples with explicit calculation done in the discrete Wigner and SIC-POVM representation — the two most widely used frame representations for discrete quantum mechanics. We further comment on the non-classicality of the evolution and believe that the process’ generator, being the one possessing the negativity, should be a focus for future work in its role in quantum non-Markovianity.
@@
Acknowledgements.—We thank Dagomir Kaszlikowski and Lim Zi Yao for discussions. This project was supported by the Ministry of Education, Singapore, under its Academic Research Fund Programme (RG154/24). K.O. was supported by grants from the National Research Foundation, Singapore and A*STAR under its Quantum Engineering Programme 2.0 (NRF2021-QEP2-02-P07) and the Singapore Ministry of Education Academic Research Fund (T2EP50222-0038).
Appendix: Discrete Wigner representation for qutrit—The following are the frame operators for the discrete Wigner representation in $d=3$:
$F_{00}=\frac{1}{3}\begin{bmatrix}1&0&0\\ 0&0&1\\ 0&1&0\end{bmatrix}\,,\quad$ $F_{01}=\frac{1}{3}\begin{bmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{bmatrix}\,,$ (27)
$F_{10}=\frac{1}{3}\begin{bmatrix}0&1&0\\ 1&0&0\\ 0&0&1\end{bmatrix}\,,\quad$ $F_{10}=\frac{1}{3}\begin{bmatrix}1&0&0\\ 0&0&-\omega\\ 0&\omega^{2}&0\end{bmatrix}\,,$
$F_{11}=\frac{1}{3}\begin{bmatrix}0&0&\omega^{2}\\ 0&1&0\\ -\omega&0&0\end{bmatrix}\,,\quad$ $F_{12}=\frac{1}{3}\begin{bmatrix}0&-\omega&0\\ \omega^{2}&0&0\\ 0&0&1\end{bmatrix}\,,$
$F_{20}=\frac{1}{3}\begin{bmatrix}1&0&0\\ 0&0&\omega^{2}\\ 0&-\omega&0\end{bmatrix}\,,\quad$ $F_{21}=\frac{1}{3}\begin{bmatrix}0&0&-\omega\\ 0&1&0\\ \omega^{2}&0&0\end{bmatrix}\,,$
$F_{22}=\frac{1}{3}\begin{bmatrix}0&\omega^{2}&0\\ -\omega&0&0\\ 0&0&1\end{bmatrix}\,,\quad$

### v2

@@ V Examples
where $j\equiv(j_{1},j_{2})\in\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ and $\sigma_{x}$, $\sigma_{y}$, and $\sigma_{z}$ are qubit Pauli operators. The construction for Gross’s discrete Wigner representation is more involved and since we are only working with a qutrit system in the example below, we show the frame operators for $d=3$ in the Appendix.
We ascertain that the results below are frame independent by repeating the calculation with another representation, namely, the symmetric informationally complete (SIC) POVM representation [37].
V.1 Pure decoherence model
Let us start with the prototypical model of a two-level system in a bosonic bath environment undergoing pure decoherence dynamics [11]. The evolution of the system’s state is given by the master equation
@@ V.1 Pure decoherence model
where the time-dependent decay rate $\gamma(t)=-\frac{1}{G(t)}\frac{d}{dt}G(t)$, with $G(t)$ the decoherence function. From this the dynamical map reads
@@
Note that the map $\Lambda_{t}$ is completely positive and trace preserving for $G(t)\leq 1$.
To find the non-Markovianity criterion, we first find the quasistochastic representation of $\Lambda_{t}$ following the Wootters representation:
@@
where $G\equiv G(t)$ for simplicity. Notice that here $S^{\Lambda_{t}}$ is self-transpose; hence $\Lambda_{t}$ is its self-dual, i.e., $\Lambda_{t}^{\dagger}=\Lambda_{t}$. More importantly, the quasibistochastic map is represented non-negatively (same as the SIC POVM representation), which allows us to use the main result in Eq. (9). In fact, all the examples in this paper have a map that is non-negative. Calculating the left hand side of Eq. (11), we find its eigenvalues are given by $\{1,1,G^{2},G^{2}\}$. Using Eq. (IV), we obtain that
@@
Therefore, it is clear that the Markovian condition is violated when $\gamma(t)<0$. This criterion coincides with the BLP measure and in fact with the P- and CP-divisibility measures [14]. In contrast, the non-Markovianity measures are different in general between the measure introduced here and the BLP measure, which read $\mathcal{N}=-4\int_{\gamma<0}dt\,\gamma(t)G^{2}(t)$ and $\mathcal{N}_{\text{BLP}}=-2\int_{\gamma<0}dt\,\gamma(t)G(t)$, respectively.
V.2 Dissipation model
Next we consider the same system and bath environment but with a system-environment interaction that allows for dissipation in the populations too. This evolution is governed by the time-local master equation
@@ V.2 Dissipation model
$\Lambda_{t}:\rho(0)\rightarrow\rho(t)=\begin{bmatrix}\rho_{00}+[1-|G(t)|^{2}]\rho_{11}&G(t)^{*}\rho_{01}\\ G(t)\rho_{10}&|G(t)|^{2}\rho_{11}\end{bmatrix}\,$. (19)
@@
$\zeta(t)=8|G(t)|^{3}\frac{d}{dt}|G(t)|=-4\gamma(t)|G(t)|^{4}$ (20)
and identify that non-Markovianity is observed when $\gamma(t)<0$. Again, the witness here coincides with the BLP and CP-divisibility measures [1].
V.3 Random unitary channel
In the last example, we look into evolution under a random unitary channel of a qubit and a qutrit system. The general qudit system has been studied in [38]. Here we consider the master equation
@@ V.3 Random unitary channel
with $\Gamma_{k}(t)=\int_{0}^{t}\gamma_{k}(\tau)d\tau$ and $H_{\alpha\beta}$ a complex Hadamard matrix $H_{\alpha\beta}=\omega^{\alpha\times\beta}$, $\alpha\times\beta\equiv(k,l)\times(m,n)=kn-lm$. Note that for $p_{\alpha}(t)$ to be a probability distribution at all times and normalized, $\sum_{\alpha=0}^{d^{2}-1}p_{\alpha}(t)=1$, we set $\lambda_{0}(t)=1$. From the above relations, we also know that there are $d^{2}-1$ independent rates $\gamma_{k}(t)$, and $\gamma_{0}$ can be determined by the other $\gamma_{k}$’s via $\gamma_{0}=-\sum_{k=1}^{d^{2}-1}\gamma_{k}$.
For $d=2$, the Weyl operators reduce to the identity and Pauli operators. In this case, we find the criteria for the Markovian condition to be
@@
This coincides with the BLP and P-divisibility conditions, as expected [39], but differs from the CP-divisibility condition, which is $\gamma_{k}(t)\geq 0$. This is an instance of the departure of the Markovian condition between information backflow and divisibility, with the former being weaker.
In $d=3$, we find the Markovian criteria as
@@
To our knowledge, characterization based on BLP and P-divisibility appears to be highly nontrivial and only a sufficient (but not necessary) condition is obtained,
@@
for all triples $\{i_{1},i_{2},i_{3}\}\subset\{1,\dots,8\}$. Similar to the qubit case, the evolution is CP-divisible if all rates are non-negative $\gamma_{\alpha}(t)\geq 0$.
VI Discussions
@@
The monotonic relation in Eq. (11) provides an interesting perspective on non-Markovianity between classical and quantum theory. In the classical case, evolution is generally described by a stochastic map. A bistochastic map, a subset of all stochastic maps, has a special property such that its transpose returns a valid map as well. One can think of the transpose as an attempt to “reverse” the transition probabilities between the input and output. Recent works [40, 41, 42] have also pointed out that the Hermitian adjoint and transpose of a quantum and a classical process, respectively, correspond to the time-reversal of the process. Interestingly, under the QPR the time-reversal notion has been unified for both theories with the transpose operation. Reversal of the process, however, does not always help in reversing the state of the system. This is also reflected in the monotonic relation, in both forms of $S^{T}S$ and $SS^{T}$, that will mix any probability distribution further. In contrast, applying forward and reverse quantum processes successively, or vice versa, on any state may result in a momentary revival of quantum features, albeit still leading to noise eventually.
@@
It should be noted that reversal is not equivalent to the inverse, which is generally not a valid map except for unitaries. [^2] For classical theory, unitary processes are described by permutation matrices, i.e., matrices with a single $1$ in each row and column and 0 elsewhere. Nevertheless, for an irreversible process one can still construct a map that approximately inverts the forward process known as Bayesian inference. Remarkably, the canonical quantum Bayesian inference, the Petz map, also contains $S^{T}$, which might imply that it is a useful ingredient in retrieving lost information [43].
@@
In formulating the condition for Markovianity, it is also required that quantum maps are represented non-negatively. This is seemingly counterintuitive to current perception, which widely believes that the appearance of negativity is necessary for many quantum phenomena. Thus, one would also think that quantum non-Markovianity, a nonclassical phenomenon, would occur due to negativity in the dynamical map. It would be interesting to show if a negative quasi-stochastic map corresponding to a Markovian evolution would still satisfy the monotonic relation in Eq. (9). Alternatively, it would also be interesting if it is always possible to find a frame representation that will return a non-negative quasistochastic matrix.
@@
Although the QPR of the dynamical maps studied above are represented non-negatively, we would like to point out that the generator $\mathcal{L}$ has a nonclassical representation. More explicitly, we find $L=[L_{jk}]$, $L_{jk}=\text{Tr}\left[F_{j}\mathcal{L}[G_{k}]\right]\in\mathbb{R}$, which is a $d^{2}\times d^{2}$ matrix that satisfies $\sum_{j}L_{jk}=0\,\forall\,k$. This is the quantum analog of the classical Kolmogorov generator [44] that additionally satisfies $L_{jk}\geq 0$ for $j\neq k$, i.e., its off-diagonal elements are non-negative. For the pure decoherence and random unitary channel models, we observe that the off-diagonal entries of $L$ become negative when the non-Markovianity condition is satisfied and non-negative otherwise. However, for the dissipative model the matrix $L$ has negative nondiagonal entries in general. This suggests that negativity in the generator appears to be necessary but not sufficient for non-Markovianity. This observation has also been pointed out in [45], which studied its connection to classical and quantum P-divisibility. It remains unclear how exactly negativity in the generator of quantum evolution contributes to the degree of non-Markovianity. We believe that the study using the QPR would be crucial in gaining a fundamental understanding of the difference between classical and quantum processes.
VII Conclusions
@@
In this work, we have proposed a new witness and measure of non-Markovianity using the formalism of quasiprobability representation. In particular, the measure detects information backflow through the breaking of the monotonic relation of the product of a quasibistochastic map, representing the dynamical map, and its transpose. The measure is explicitly dependent only on the dynamical map, thus forgoing any optimization task that plagues much of the previously proposed measure. This is particularly useful when studying the evolution of a quantum system of large dimensions. We illustrated the use over several paradigmatic examples with the explicit calculation done in the discrete Wigner and SIC POVM representation, the two most widely used frame representations for discrete quantum mechanics. We further discussed on the nonclassicality of the evolution and believe that the process generator, the one possessing the negativity, should be a focus for future work in its role in quantum non-Markovianity.
It should be noted that our study of non-Markovian dynamics is to the extent of the information backflow notion, and there are several other concepts of non-Markovianity that are not explored here. Although the concept of divisibility seems to be the popular definition of Markovianity, there remains a large debate to this day on the proper definition of Markovianity [2]. We also would like to mention that concepts beyond quantum dynamical maps, used primarily in divisibility and information backflow approach, also exist, e.g., the quantum regression formula [3, 46]. In this case, we believe that recent advances in other forms of quasiprobability, namely, Kirkwood-Dirac quasiprobability [47], can be an indispensable tool to study quantum regression formula as it concerns time-ordered correlation functions, a quantity that can be efficiently described using Kirkwood-Dirac quasiprobability [48]. We leave this for future work.
Acknowledgements
@@
We thank Dagomir Kaszlikowski and Lim Zi Yao for discussions. This work was supported by the Ministry of Education, Singapore, under its Academic Research Fund Programme (Grant No. RG154/24). K.O. was supported by grants from the National Research Foundation, Singapore and A*STAR under its Quantum Engineering Programme 2.0 (Grant No. NRF2021-QEP2-02-P07) and the Singapore Ministry of Education Academic Research Fund (Grant No. MOE-T2EP50222-0017).
Appendix A Appendix: Discrete Wigner representation for qutrit
The following are the frame operators for the discrete Wigner representation in $d=3$:
$F_{00}=\frac{1}{3}\begin{bmatrix}1&0&0\\ 0&0&1\\ 0&1&0\end{bmatrix}\,$$\quad$ $F_{01}=\frac{1}{3}\begin{bmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{bmatrix}\,$, (27)
$F_{10}=\frac{1}{3}\begin{bmatrix}0&1&0\\ 1&0&0\\ 0&0&1\end{bmatrix}\,$$\quad$ $F_{10}=\frac{1}{3}\begin{bmatrix}1&0&0\\ 0&0&-\omega\\ 0&\omega^{2}&0\end{bmatrix}\,$,
$F_{11}=\frac{1}{3}\begin{bmatrix}0&0&\omega^{2}\\ 0&1&0\\ -\omega&0&0\end{bmatrix}\,$$\quad$ $F_{12}=\frac{1}{3}\begin{bmatrix}0&-\omega&0\\ \omega^{2}&0&0\\ 0&0&1\end{bmatrix}\,$,
$F_{20}=\frac{1}{3}\begin{bmatrix}1&0&0\\ 0&0&\omega^{2}\\ 0&-\omega&0\end{bmatrix}\,$$\quad$ $F_{21}=\frac{1}{3}\begin{bmatrix}0&0&-\omega\\ 0&1&0\\ \omega^{2}&0&0\end{bmatrix}\,$,
$F_{22}=\frac{1}{3}\begin{bmatrix}0&\omega^{2}&0\\ -\omega&0&0\\ 0&0&1\end{bmatrix}\,$$\quad$